OpenStudy (anonymous):

I took a test , I don't have the answers available, and i need to know how much did i score. Topic:- Quadratic equations Would anyone help me

OpenStudy (anonymous):

If alpha and beta are the roots of the equation ax^2+bx+c=0, then the quadratic equation ax^2-bx(x-1)+c(x-1)^2=0 has roots?

OpenStudy (anonymous):

My answer:- \[\huge \frac{ \alpha }{ \alpha+1 } , \frac{ \beta }{ \beta+1 }\]

OpenStudy (anonymous):

If the roots of the equation x^2 +2cx +ab =0 are real and unequal, then the roots of the equation x^2-2(a+b)x+(a^2+b^2+2c^2) =0 are:

OpenStudy (anonymous):

My answer: imaginary

OpenStudy (anonymous):

If -3+5i is a root of the equation x^2+px+q=0 , then the ordered paire (p,q) is (p,q belong to real numbers)

OpenStudy (anonymous):

My answer: \[\huge (6,34)\]